OpenStudy (anonymous):

The half-life of a specific radioactive compound is 20 years. The equation used to model this half-life is y = 5(1/2)t/20, where t represents time. Currently, there are 5 pounds of this compound 15 feet below the earth's surface. When will there only be 4 pounds of this compound left? Round your answer to the nearest tenth of a year.

OpenStudy (mathstudent55):

Is this the equation, where t/20 is an exponent? \(y = 5 \left(\dfrac{1}{2} \right)^{\frac{t}{20}} \)

OpenStudy (anonymous):

yes i think soo

OpenStudy (anonymous):

whats t ?

OpenStudy (mathstudent55):

Let y = 4 and solve for t.

OpenStudy (anonymous):

2.5?

OpenStudy (mathstudent55):

\(4 = 5 \left(\dfrac{1}{2} \right)^{\frac{t}{20}}\) Divide both sides by 5: \(\dfrac{4}{5} = \left(\dfrac{1}{2} \right)^{\frac{t}{20}}\) Now take log of both sides.

OpenStudy (mathstudent55):

\(\log \dfrac{4}{5} = \log \left[ \left(\dfrac{1}{2} \right)^{\frac{t}{20}} \right]\)

OpenStudy (anonymous):

i divide (1/2) with what 20..?

OpenStudy (mathstudent55):

\(\log \dfrac{4}{5} = \dfrac{t}{20}\log \dfrac{1}{2} \) \(\dfrac{\log 4 - \log 5}{\log1 - \log2} = \dfrac{t}{20} \) \(\dfrac{20(\log 4 - \log 5)}{- \log2} = t \) \(\dfrac{20(\log 4 - \log 5)}{- \log2} = t \) \(\dfrac{20( \log 5 - 2\log 2) }{\log2} = t \)

OpenStudy (anonymous):

soo 5-2/2 ?

OpenStudy (anonymous):

1.5

OpenStudy (mathstudent55):

t = 6.4 years

OpenStudy (anonymous):

is this rounded ?

OpenStudy (mathstudent55):

Yes, it's rounded to 1 decimal place.